OpenStudy (emogirlatemoooocow):

Sara had 7/8 pound of turkey. She used 2/3 pound for sandwiches. How many pounds of turkey does Sara have left? A. 5/25 pound B. 1/4 pound C. 5/12 pound D. 5/8 pound

OpenStudy (knov):

hint: multiply it to find out the left pounds.

OpenStudy (emogirlatemoooocow):

@knov i got \[\frac{ 7 }{ 12}\]

OpenStudy (knov):

that's an amount of pound sara has left.

OpenStudy (emogirlatemoooocow):

but thats not a answer on there But i subtracted \[\frac{ 7 }{ 8 } and \frac{ 2 }{ 3 }\]

OpenStudy (emogirlatemoooocow):

and got \[\frac{ 5 }{ 25 }\]

OpenStudy (emogirlatemoooocow):

and got \[\frac{ 5 }{ 25 }\]

OpenStudy (knov):

She had 7/8 pounds turkey. And she USED 2/3 of it. Which means